### 28 Apr 2026 v4

# Suiren-1.0 Technical Report: A Family of Molecular Foundation Models

Abstract
We introduce Suiren-1.0, a family of molecular foundation models for the accurate modeling of diverse organic systems. Suiren-1.0 comprising three specialized variants (Suiren-Base, Suiren-Dimer, and Suiren-ConfAvg) is integrated within an algorithmic framework that bridges the gap between 3D conformational geometry and 2D statistical ensemble spaces. We first pre-train Suiren-Base (1.8B parameters) on a 70M-sample Density Functional Theory dataset using spatial self-supervision and SE(3)-equivariant architectures, achieving robust performance in quantum property prediction. Suiren-Dimer extends this capability through continued pre-training on 13.5M intermolecular interaction samples. To enable efficient downstream application, we propose Conformation Compression Distillation (CCD), a diffusion-based framework that distills complex 3D structural representations into 2D conformation-averaged representations. This yields the lightweight Suiren-ConfAvg, which generates high-fidelity representations from SMILES or molecular graphs. Our extensive evaluations demonstrate that Suiren-1.0 establishes state-of-the-art results across a range of tasks. All models and benchmarks are open-sourced.
Model Links and Resources
Figure 1: Benchmark performance of Suiren-1.0 and its counterparts. We use the normalized MAE scores ($\uparrow$).
(a) Critical & Saturation Properties
(b) Energetic Properties
(c) Fluctuation Properties
(d) Safety Properties
(e) Solution Properties
(f) Structural Properties
(g) Thermal Properties
(h) Transport Properties
Figure 2: Comparison of Suiren-1.0 model and molecular Foundation model across various tasks in 8 domains. All tasks are regression tasks, with MAE ($\downarrow$) as the evaluation metric. Due to significant differences in metric ranges across different tasks, the y-axis is scaled.
1 Introduction
Foundation models have catalyzed a paradigm shift in natural language processing and computer vision, where large-scale pre-training facilitates robust transferability across diverse downstream tasks [achiam2023gpt, team2023gemini, yang2025qwen3, liu2024deepseek]. In the science domain, pioneering molecular architectures such as MoleBERT [xia2023mole], Uni-Mol [zhou2023uni, ji2024uni], and UMA [wood2025family] have demonstrated significant promise. However, compared to the linguistic and visual domains, universal molecular modeling remains hindered by inherent scientific complexities and a scarcity of high-quality supervised data. We identify the primary challenges as follows:
  • First, the "physical priors" governing molecular systems are exceptionally complex. Molecular behavior is dictated by the intricate laws, such as quantum mechanics (e.g., the Schrödinger equation) and statistical thermodynamics (e.g., Boltzmann distributions) [schleich2013schrodinger, charbonneau1982linear]. Capturing these fundamental mechanisms solely through data-driven learning is challenging, particularly given the sparsity of high-fidelity labeled data.
  • Second, a persistent multiscale gap remains between microscopic structures and macroscopic observables. Microscopic tasks typically demand the resolution of explicit 3D conformations and electronic densities, where Density Functional Theory (DFT) enables the generation of abundant, high-quality labeled data [liu2024open, chanussot2021open, levine2025open]. Conversely, macroscopic tasks often rely on 1D SMILES or 2D molecular graphs that lack explicit conformational information. While these tasks span broad chemical spaces, their data is often scarce, as macroscopic labels frequently require costly wet-lab experiments or molecular dynamics simulations. Physically, these two modalities are intrinsically linked: macroscopic features emerge from the ensemble-averaged properties of a molecule’s conformations, governed by the Boltzmann distribution (see Figure 3). However, existing approaches largely fail to bridge this divide. Pure 3D foundation models, such as UMA, learn rich 3D representations from labeled data but lack generalizability across broad chemical tasks; meanwhile, pure 2D models, such as Mole-BERT, capture graph topology through self-supervised learning yet remain "conformation-blind," limiting their predictive effectiveness.
Figure 3: Microscopic and macroscopic representations of molecular ensembles. (a) Molecular Representation: A single molecular identity corresponds to a diverse ensemble of 3D conformations at the microscopic space. (b) Conformational distribution: The relative probability of these conformations is governed by the Boltzmann distribution as a function of potential energy. (c) Ensemble Property: Macroscopic observables emerge as the ensemble-averaged properties derived from the collective contributions of all constituent conformations.
Suiren-1.0 is designed to bridge the multiscale gap between microscopic and macroscopic representations. We first pre-train Suiren-Base (1.8B parameters) on large-scale, first-principles quantum chemical data using objectives specifically tailored for microscopic, conformation-aware representation learning. We then introduce Conformation Compression Distillation (CCD), a diffusion-based strategy that distills the knowledge of Suiren-Base into Suiren-ConfAvg. This process encodes a macroscopic latent representation that can be inverted into specific 3D conformations through energy-conditioned queries. In the absence of an energy query, Suiren-ConfAvg accepts 2D molecular graphs or 1D SMILES as input to produce generalizable molecular embeddings suitable for a wide range of downstream tasks, including materials discovery, drug design, and battery chemistry.
We evaluate Suiren-1.0 across a comprehensive suite of over 50+ tasks spanning 9 diverse scientific domains. To ensure a rigorous and fair assessment, we eschew task-specific engineering in favor of a unified fine-tuning and inference pipeline across all benchmarks. As illustrated in Figure 1 and Figure 2, Suiren-1.0 achieves consistent, state-of-the-art (SOTA) performance in the vast majority of cases. Notably, Suiren-1.0 delivers performance gains exceeding 20% on more than 20+ tasks compared to existing models. We attribute this success to the synergy between large-scale model scaling and the principled integration of physical priors.
Our main contributions are as follows:
Modeling Framework: Microscopic–Macroscopic Bridging
  • We establish a three-stage framework to unify molecular scales: (i) pre-training a 3D conformation-aware foundation model for high-fidelity microscopic representation learning; (ii) distilling this knowledge into a compressed, conformation-agnostic model for macroscopic adaptation via Conformation Compression Distillation; and (iii) fine-tuning task-specific encoders for a diverse suite of downstream scientific applications.
Pre-Training: Physical Priors and First-Principles Data
  • We train Suiren-Base on large-scale first-principles quantum-chemical data (Qo2mol liu2024open) and incorporate physically motivated algorithms, including advanced EMPP [an2025empp] and advanced EST [an2025est], to improve representation quality.
  • The resulting representations capture conformation-sensitive microscopic information and transfer effectively across downstream scientific tasks. We further provide a continued pre-training variant for dimer systems (Suiren-Dimer).
Transfer Learning: Broad Molecular Applicability
  • By distilling from Suiren-Base to Suiren-ConfAvg, we enable strong performance when only graph or SMILES inputs are available, improving deployability in real-world molecular pipelines.
  • We benchmark Suiren-ConfAvg on 50+ property prediction tasks and observe robust improvements over advanced molecular baselines.
Open Science
  • We provide a description of pre-training, distillation, and finetune models, and release their weights to facilitate reproducible molecular foundation-model research. Furthermore, we open a comprehensive benchmark MoleHB for molecular model evaluation.
The remainder of this paper is organized as follows. Section 2 details the architectural design and the broader Suiren-1.0 model family. Section 3 describes our data curation and pre-training methodology, followed by Section 4, which introduces our post-training distillation and fine-tuning protocols. In Section 5, we present a comprehensive evaluation of Suiren-1.0. Finally, Section 6 concludes the paper with a discussion on current limitations and future research directions.
2 Architecture
2.1 Large SO(3)-Equivariant Graph Neural Network
Suiren-Base is a high-degree equivariant graph neural network (GNN) designed for 3D conformational representation learning. The architecture integrates an EquiformerV2 model [liao2023equiformerv2] with a dense Mixture-of-Experts (MoE) update block, which concurrently utilizes both S2Activation and Equivariant Spherical Transformer (EST) experts [an2025est] in each forward pass. Following the standard message-passing framework [gilmer2017neural], the model architecture is formulated as:
$\mathbf{m}_{i}^{(k)}$ $=\sum_{j\in\mathcal{N}(i)}\psi_{m}^{(l)}\left(\mathbf{x}_{i}^{(l)},\mathbf{x}_{j}^{(l)},\mathbf{e}_{ij}\right),$ (1)
$\mathbf{x}_{i}^{(k+1)}$ $=\psi_{u}^{(l)}\left(\mathbf{x}_{i}^{(l)},\mathbf{m}_{i}^{(l)}\right),$ (2)
where $N(i)$ denotes the neighbor set of node $i$, $\mathbf{x}_{i}$ represents the node embeddings, $\mathbf{e}_{ij}$ denotes the edge features, and $\psi_{m}^{(l)}$ and $\psi_{u}^{(l)}$ correspond to the message and update functions, respectively. The message block captures interatomic interactions with a computational complexity linear in the number of edges, and aggregates these into node-level messages. Conversely, the update block processes these messages with a complexity linear in the number of nodes. These components serve functional roles analogous to the self-attention and feed-forward network (FFN) modules in standard Transformer architectures.
Figure 4 illustrates the architecture of Suiren-Base. The message block is adapted from the EquiformerV2 graph attention block and utilizes an $SO(2)$-linear operation to integrate edge features $\mathbf{e}_{ij}$ with node attributes $(\mathbf{x}_{i},\mathbf{x}_{j})$. For the update block, we employ a dense MoE design to enhance model capacity. This architectural choice is driven by two primary observations: (1) the update block is computationally efficient, allowing the addition of experts with minimal latency overhead; and (2) the complexity of the aggregated messages benefits significantly from increased parameter capacity.
Suiren-Base contains 20 layers ($K=20$), and each MoE block contains 20 S2Activation experts and 20 EST experts, balancing equivariance and expressiveness. The original EST maps steerable group embeddings from the harmonic domain to spatial domain with Fourier basis sampled spherical points, updates embeddings with a spherical Transformer, and projects them back:
$f(\vec{\mathbf{p}})$ $=\mathcal{F}(\mathbf{x})=\sum_{l=0}^{\infty}\sum_{m=-l}^{l}\mathbf{x}^{(l,m)}Y^{(l,m)}(\vec{\mathbf{p}})$ # Fourier transform on a single spherical point (3)
$\mathbf{f}$ $=[f(\vec{\mathbf{p}}_{1}),f(\vec{\mathbf{p}}_{2}),...,f(\vec{\mathbf{p}}_{S})]$ # The spatial representation on sampling points $\vec{\mathbf{p}}_{i}$ (4)
$\hat{\mathbf{f}}$ $=\mathrm{Trans}([\mathbf{f};\mathbf{P}])$ # Transformer with orientation embedding $\mathbf{P}=[\vec{\mathbf{p}}_{1},...,\vec{\mathbf{p}}_{S}]$ (5)
$\hat{\mathbf{x}}$ $=\sum_{s=1}^{S}\hat{\mathbf{f}}_{s}\cdot\mathbf{Y}^{*}(\vec{\mathbf{p_{s}}})$ # Back to harmonic domain (6)
Figure 4: The architecture of the Suiren-Base model. (a) Overall framework. (b) A dense MoE block. (c) Modified EST expert: during training, the spherical Fourier transform basis set and orientation embedding are subjected to a random rotation.
where $Y^{l,m}(\cdot)$ and $\mathbf{Y}(\cdot)=[Y^{l_{1},m_{1}}(\cdot),Y^{l_{2},m_{2}}(\cdot),$...] denote the spherical harmonic basis (scale and vector), $\vec{\mathbf{p}}$ denotes the sample point (orientation) in the sphere, and the term $\hat{\cdot}$ is used to represent the embedding update. Although uniform spherical sampling in EST offers partial equivariance, it remains susceptible to discretization-induced errors. To mitigate these artifacts, we propose a basis-rotation strategy for EST experts. During training, the Fourier basis for each sample in a mini-batch is pre-rotated by a random 3D rotation $\mathbf{R}$. Since this procedure only modifies the basis orientation, the computational overhead remains negligible. This approach exposes the model to a diverse range of orientations, encouraging the learning of orientation-consistent responses and more closely approximating continuous spherical Fourier behavior. The formulation is defined as:
$\mathbf{f}$ $=[f(\mathbf{R}\vec{\mathbf{p}}_{1}),f(\mathbf{R}\vec{\mathbf{p}}_{2}),...,f(\mathbf{R}\vec{\mathbf{p}}_{S})]$ (7)
$\hat{\mathbf{f}}$ $=\mathrm{Trans}([\mathbf{f};\mathbf{R}\mathbf{P}])$ (8)
$\hat{\mathbf{x}}$ $=\sum_{s=1}^{S}\hat{\mathbf{f}}_{s}\cdot\mathbf{Y}^{*}(\mathbf{R}\vec{\mathbf{p_{s}}}).$ (9)
By leveraging this adaptive equivariance mechanism, the spherical sampling density $S$ can be optimized toward the Nyquist-rate lower bound, $S\geqslant(2L)^{2}$, where $L$ denotes the maximum degree of the spherical harmonic embedding. This reduction significantly lowers both training and inference overhead without compromising the robustness of the model’s equivariant properties.
Both Suiren-Base and Suiren-Dimer utilize a unified backbone architecture. In practice, these models take 3D atomic coordinates of molecules or dimers as input to predict quantum-accurate (DFT-level) potential energies and interatomic forces.
2.2 Conformation Compression Distillation
As illustrated in Figure 3, molecular properties are typically determined by ensemble-averaged behavior across multiple physically plausible conformers. While these conformer probabilities are governed by the Boltzmann distribution, the exact distribution—and the underlying potential energy surface (PES)—is generally unknown a priori. To address this, we propose Conformation Compression Distillation (CCD), a feature-alignment framework designed for one-to-many molecule-conformer mapping.
For each molecule-conformer pair, CCD operates on two distinct modalities: the 2D molecular topology (SMILES or graph) and the 3D conformer with its associated energy $E$. The 2D input is processed via a Graph Attention Network (GAT) to extract a latent representation $\mathbf{h}^{\mathrm{2D}}$, while the 3D conformer is encoded by a pre-trained Suiren-Base teacher to yield an equivariant representation $\mathbf{h}^{\mathrm{3D}}$. We then introduce a 3D diffusion-based model featuring a lightweight Equiformer+MoE+EST dynamics network $\varphi_{\theta}(\cdot)$. This network is conditioned on both $\mathbf{h}^{\mathrm{2D}}$ and the energy $E$, with the diffusion process targeting the joint reconstruction of the 3D representation $\mathbf{h}^{\mathrm{3D}}$ and the molecular 3D coordinates $\mathbf{c}$ (see Figure 5(b)).
During training, we add random noise to the clean 3D target state:
$\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{-1}+\sigma_{t}\boldsymbol{\varepsilon}$ (10)
where $\mathbf{z}_{-1}=[\mathbf{h}^{\mathrm{3D}};\mathbf{r}]$ denotes the clean target state, $\mathbf{z}_{t}$ is the noisy state at timestep $t\in[0,T]$, $\alpha_{t},\sigma_{t}$ are schedule coefficients, and $\boldsymbol{\varepsilon}$ denotes the Gaussian noise. We freeze the weights of Suiren-Base, and train the dynamics network and 2D representation model by predicting the noise:
$\hat{\boldsymbol{\varepsilon}}$ $=\varphi_{\theta}(\mathbf{z}_{t},t,\mathbf{h}^{\mathrm{2D}},E)$ (11)
$\mathcal{L}_{\mathrm{CCD}}$ $=\operatorname{MSE}\!\left(\hat{\boldsymbol{\varepsilon}},\boldsymbol{\varepsilon}\right),$ (12)
where $E$ is encoded with Gaussian embeddings and $\mathcal{L}_{\mathrm{CCD}}$ denote optimize objective functions.
Upon completion of training, the framework yields two primary components: a 2D molecular encoder and a generative diffusion dynamics model. These modules facilitate both high-fidelity conformer generation and robust 2D representation learning. For conformer generation, the diffusion dynamics model captures the multimodal nature of the structural space, generating diverse ensembles rather than collapsing to a single low-energy mode [Landrum2016RDKit2016_09_4, xu2024gtmgc]. The reverse-time sampling step from timestep $t$ to s=t1s=t-1 is formulated as:
$\mathbf{z}_{s}=\frac{1}{\alpha_{t|s}}\mathbf{z}_{t}-\frac{\sigma_{t|s}^{2}}{\alpha_{t|s}\sigma_{t}}\cdot\varphi_{\theta}(\mathbf{z}_{t},t,\mathbf{h}^{\mathrm{2D}},E)+\sigma_{t\to s}\cdot\boldsymbol{\varepsilon},$ (13)
where $\alpha_{t|s}=\alpha_{t}/\alpha_{s}$, $\sigma_{t|s}^{2}=\sigma_{t}^{2}-\alpha_{t|s}^{2}\sigma_{s}^{2}$, and $\sigma_{t\to s}=\frac{\sigma_{t|s}\sigma_{s}}{\sigma_{t}}$. Regarding 2D representation learning, CCD implicitly characterizes the mapping from 2D graphs to 3D configurations. Given the significant modality gap between these spaces, direct feature alignment is often ill-posed. The diffusion strategy in CCD addresses this by enabling the 2D encoder to reconstruct 3D information in stages, thereby mitigating optimization challenges. This process yields the Suiren-ConfAvg model, which provides versatile representations for a broad range of macroscopic molecular tasks.
Figure 5: Overview of training stages. (a) 3D Pre-training: Self-supervised learning on 3D molecular conformations. (b) Conformation Distillation: Distilling 3D geometric knowledge into a conformation-averaged representation. (c) Downstream Fine-tuning: Adapting the model for supervised molecular property prediction.
2.3 Dual Graph Neural Network
Following the CCD-based training of the 2D representation model, we propose a Dual Graph Neural Network (DGNN) architecture for downstream fine-tuning. As illustrated in Figure 5(c), the DGNN consists of two parallel sub-networks: a pre-trained Suiren-ConfAvg module, initialized via CCD, and a randomly initialized task-specific GNN. During the forward pass, latent representations from Suiren-ConfAvg are injected into the corresponding layers of the task-specific GNN to provide structural guidance. To mitigate catastrophic forgetting and preserve the learned conformation-averaged features, the Suiren-ConfAvg weights remain frozen throughout this stage. While both modules utilize GAT architectures, the task-specific GNN is designed with greater depth to absorb all representations of Suiren-ConfAvg.
3 Pre-training
In this section, we describe the construction of the pre-training data, the multi-stage training pipeline, and the evaluation of the resulting base models.
3.1 Pre-training Data
The Suiren-1.0 model utilizes a vast corpus of first-principles molecular data for pre-training. Using Density Functional Theory (DFT) at the B3LYP/def2-SVP level, we generated 70 million conformer samples for organic molecules encompassing the H, C, N, O, F, P, S, Cl, Br, and I elements. Of these, 20 million samples have been publicly released as the Qo2mol dataset [liu2024open]. Each entry includes Cartesian coordinates, energies, forces, trajectory information, and associated metadata. Prior to training, we perform rigorous data cleaning to remove anomalous samples and identify the terminal optimized geometry of each trajectory, which serves as an auxiliary supervision target.
To enhance data efficiency, we augment the training process using EMPP method [an2025empp]. For each molecule, a random atom is deleted rather than masked, and the model is required to reconstruct its coordinates conditioned on the atom type and target molecular energy. This objective encourages the model to learn physically plausible local potential-energy landscapes and effectively doubles the training volume. We further refine the original EMPP formulation: rather than employing layer-wise conditioning of the deleted-atom signals, we feed these inputs exclusively to an EMPP-specific coordinate-prediction head. This modification ensures that the shared backbone maintains a consistent forward pass across all pre-training objectives.
3.2 Pre-training Stage
Pre-training is divided into three stages.
Stage 1 (multi-task foundational capability learning)
Within the Fairchem framework, we train a 1.8B-parameter 3D model using both the original dataset and the EMPP-augmented samples. The pre-training tasks include energy prediction, force prediction, optimized-trajectory endpoint structure prediction, optimized-trajectory endpoint energy prediction, and EMPP missing-coordinate completion. All task losses are optimized jointly. Inspired by curriculum learning, we prioritize smaller molecular systems (fewer atoms) in earlier training phases. The weights of endpoint-structure prediction and endpoint-energy prediction losses are also gradually increased during training.
Stage 1 is trained on 320 NVIDIA H800 GPUs with mixed precision and graph parallelization. Because PyTorch Geometric supports a variable number of atoms per mini-batch, we combine dynamic batch balancing and activation recomputation to avoid memory overflow.
Stage 2 (core capability refinement)
For regression targets such as energy and force, mixed-precision training can substantially degrade model performance. Nonetheless, this degradation can be corrected by a relatively short fine-tuning stage. In Stage 2, we use only the 70M original samples for full-precision fine-tuning, with the following tasks: energy prediction, force prediction, optimized-trajectory endpoint structure prediction, and optimized-trajectory endpoint energy prediction. The weights for endpoint-structure prediction and endpoint-energy prediction are fixed to a small constant. Except for the switch from mixed precision to full precision, all optimization strategies remain unchanged.
After extensive hyperparameter search across the first two stages, we obtain Suiren-Base.
Stage 3 (continued pre-training in the dimer domain)
Suiren-Base primarily learns intra-molecular interactions. For applications such as drug design, inter-molecular interactions, including long-range impacts, are often essential. To address this, we generate 13.5M dimer samples with DFT and continue pre-training from Suiren-Base. The architecture and optimization recipe remain consistent with Stage 2, yielding the dimer-focused model Suiren-Dimer.
3.3 Pre-training Evaluation
We evaluate pre-training quality primarily with MAE on energy prediction, force prediction, and optimized-trajectory endpoint prediction. For investigate pre-training performance, we reproduce several strong baselines on a Qo2mol subset. We also include UMA-family results as an external performance anchor and analyze MoE routing statistics.
Standard Evaluation
We sample validation subsets from Qo2mol containing more than 1M conformers across different molecular scales. Suiren-Base and Suiren-Dimer use different subsets. As shown in Table 1, Suiren-Base achieves highly accurate results on both energy and force prediction. Optimized-trajectory endpoint structure and endpoint energy are substantially more challenging targets, yet Suiren-1.0 still attains strong accuracy. Note that Suiren-0.0 is an internal transitional model trained with less parameters and a weakened backbone architecture. It uses the same training set and hyperparameters as Suiren-Base (1.0). We attribute Suiren-Base’s improvements to increased parameters, optimized backbone architecture, and refined training algorithms. Additionally, we evaluate the continued pre-training model, Suiren-Dimer. Compared with intra-molecular settings, inter-molecular trajectories are more complex. Accordingly, Suiren-Dimer is weaker than Suiren-Base on endpoint structure and endpoint energy prediction, but it still maintains strong performance on energy and force prediction. The comparison between Suiren-Base and Suiren-Dimer is indirect, as they used different validation sets.
Table 1: Comparison among energy/forces prediction models.
Suiren-0.0 Suiren-Base (1.0) Suiren-Dimer (1.0) # Total Params 1.5B 1.8B 1.8B Molecular Energy MAE (meV) $\downarrow$ 10.82 9.08 10.30 Atomic Energy MAE (meV) $\downarrow$ 0.307 0.258 0.292 Atomic Forces MAE (meV / Å) $\downarrow$ 1.472 0.770 1.294 Optimized Structure MAE (Å) $\downarrow$ 0.02895 0.02467 0.06830 Optimized Energy MAE (meV) $\downarrow$ 28.187 16.886 58.325
Feature Evaluation
We monitor MoE routing-weight distributions and observe that they become progressively sparse during training. Most mass eventually concentrates on a small subset of experts, while unused experts still retain non-negligible weights. For this reason, we do not adopt top-$K$ routing in Suiren-Base. In addition, the two kind of experts in Suiren-Base receive broadly similar aggregate routing mass, with EST experts being slightly higher on average than standard experts.
4 Post-training
4.1 Post-training Stage
Stage 1 (diffusion distillation)
We post-train on the same 70M molecules used in pre-training, but with a different objective. Follow Section 2.2, We condition a diffusion model on 2D representations and embedding of conformer energies. The diffusion model and 2D GNN learn to generate corresponding 3D representations and 3D coordinates. During this stage, the 3D branch is instantiated with Suiren-Base and kept frozen.
Stage 2 (contrastive learning)
After Stage 1 reaches a stable regime, we introduce an additional alignment objective. Specifically, we attach one projection head to the 2D model and one to the 3D model, and apply SigLIP-style contrastive learning [zhai2023sigmoid] on their outputs. The 3D branch remains frozen, and the Stage-1 diffusion objective is retained. The diffusion and contrastive objectives are jointly optimized with task-specific loss weights.
Through the first two stages, we obtain the Suiren-ConfAvg model.
Stage 3 (property prediction)
In the final stage, we evaluate the performance of Suiren-ConfAvg across a diverse array of downstream benchmarks. Each task involves predicting a specific molecular property using sparse experimental wet-lab data. We fine-tune the model for each objective using the integrated DGNN+Suiren-ConfAvg architecture. To demonstrate the general transferability and robustness of Suiren-ConfAvg representations, we maintain a unified hyperparameter configuration across all tasks, regardless of the domain. A detailed account of these configurations and the comprehensive evaluation results are documented in Section 5.
5 Experiments
5.1 Benchmark
5.1.1 MoleHB
Dataset
We introduce Molelecular handbook (MoleHB), a comprehensive molecular property prediction benchmark encompassing 40+ heterogeneous tasks. The benchmark spans several critical scientific domains, including safety, structural, critical and saturation, energetic, thermal, solution, transport and fluctuation properties. All data points are inferred from canonical handbook [yaws1999chemical] and have been rigorously validated via wet-lab experiments to ensure high-fidelity, stable values. We propose two evaluation protocols: (1) Random split: a standard random split to evaluate performance under similar data distributions; and (2) Size-Stratified split: a strategy where molecules with larger atom counts are assigned to the validation set to assess the model’s structural extrapolation capabilities. To facilitate reproducible research, we provide the MoleHB dataset and two splitting protocols on https://modelscope.cn/datasets/ajy112/MoleHB
The main paper primarily considers results from the Random split. Results of Size-Stratified split are shown in Appendix C.
Baselines and Configurations
We benchmarked three state-of-the-art models on MoleHB: MoleBERT [xia2023mole], Uni-Molv1 [zhou2023uni], and Uni-Molv2 [ji2024uni]. Like the Suiren family, these baselines utilize large-scale pre-training to generate high-quality representations for diverse molecular tasks. To ensure a fair comparison, all baselines were reproduced using their official training scripts and identical hyperparameter configurations to the Suiren models.
Detailed training configurations for all evaluated methods are summarized in Table 2. We ensured that each model reached performance convergence under these settings. All experiments were performed on a single NVIDIA RTX 4090 GPU, with fine-tuning typically completed in less than one hour per task.
Table 2: Training configurations of the Suiren model and other baselines on MoleHB experiments.
Training Hyperparameter Value
Task Formulation
Loss function MAE
Train/validation split ratio (Random and Size-Stratified) 0.8:0.2
Optimization
Mini-batch size per GPU 8
Optimizer adamw
Initial learning rate 4e-4
Weight decay coefficient 0.01
Momentum coefficient 0.9
Schedule
Scheduler type cosine
Warmup epochs 0
Minimum learning rate 1e-6
Total training epochs 200
5.1.2 Therapeutics Data Commons
Dataset
Therapeutics Data Commons (TDC) [huang2021therapeutics] is an open-access platform providing AI-ready datasets and benchmarks for drug discovery. It covers diverse therapeutic tasks—including target discovery, activity screening, efficacy, and safety—across small molecules, antibodies, and vaccines. We evaluate Suiren model on its ADMET group.
Baselines and Configurations
TDC is a public leaderboard where readers can find scores for various methods on its official website. Here, we follow [gao2023uni], using ChemProp [stokes2020deep], DeepAutoQSAR [dixon2016autoqsar], DeepPurpose [huang2020deeppurpose], and Uni-QSAR [gao2023uni] as baselines. Note that TDC ADMET tasks include both regression and classification tasks. Regression tasks use configurations identical to those in Table 2. For classification tasks, the loss function is changed to cross-entropy while all other settings remain consistent.
5.2 Results
5.2.1 MoleHB (Random Split)
Overall Performance Summary
We comprehensively evaluate the predictive performance of Suiren-ConfAvg against three representative baseline models (Mole-BERT, Uni-Mol v1, and Uni-Mol v2) across 40+ molecular properties spanning eight categories: critical & saturation, safety, fluctuation, solution, thermal, structural, energetic, and transport properties. Performance is measured using Mean Absolute Error (MAE, lower is better) and coefficient of determination (R², higher is better). As summarized in Tables 311, Suiren-ConfAvg achieves state-of-the-art MAE on 41 out of 43 properties, with consistent improvements in R² for the majority of tasks.
Table 3: Results of critical and saturation properties: model performance (MAE/R2). Best MAE and best R2 per property are boldfaced.
Property Unit Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
critical temperature K 27.3807/0.8340 9.3277/0.9787 30.6705/0.8850 6.6725/0.9803 28.47
critical pressure bar 2.0733/0.9045 0.7332/0.9765 2.9769/0.9084 0.4953/0.9804 32.45
critical density g/ml 0.0134/0.9253 0.0054/0.9814 0.0115/0.9497 0.0043/0.9782 20.74
critical volume cm3/mol 61.8034/0.8484 8.6329/0.9981 7.0201/0.9989 3.5884/0.9994 48.88
critical compressibility cm3 0.0189/0.6178 0.0066/0.8858 0.0098/0.8808 0.0053/0.9127 19.70
Table 4: Results of safety properties.
Property Unit Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
flash point F 19.6978/0.8017 9.9274/0.9472 13.9535/0.9438 8.2072/0.9645 17.33
lower explosive limit vol% 0.1689/0.6854 0.0974/0.8926 0.1106/0.8341 0.0871/0.8599 10.59
upper explosive limit vol% 1.4188/0.5846 1.0460/0.6979 1.0023/0.6942 0.7811/0.8423 22.07
Table 5: Results of fluctuation properties.
Property Unit Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
heat capacity of liquid J/mol/K 16.0687/0.7922 5.9321/0.9812 5.8939/0.9792 4.2184/0.9861 28.43
heat capacity of solid J/mol/K 49.2807/0.9499 6.5306/0.9988 3.2724/0.9990 1.0669/0.9997 67.40
coefficient of thermal expansion of liquid 1/K 0.0002/0.1716 0.0001/0.8158 0.0001/0.6211 0.0000/0.9256 60.00
heat capacity of gas J/mol/K 24.1482/0.8849 4.4112/0.9685 3.2478/0.9691 2.6712/0.9690 17.75
Table 6: Results of solution properties.
Property Unit Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
octanol water partition coefficient 0.3786/0.9414 0.1491/0.9877 0.1386/0.9858 0.1203/0.9897 13.20
solubility in water ppm(wt) 0.2110/0.9422 0.0564/0.9973 0.0456/0.9980 0.0387/0.9972 15.07
solubility in water containing salt ppm(wt) 0.2535/0.7540 0.0615/0.9924 0.0406/0.9961 0.0447/0.9931 -10.07
solubility parameter (J/cm3)1/2 0.7111/0.8331 0.3873/0.9236 0.7614/0.8282 0.3368/0.9300 13.04
henrys law constant for compound in water 0.2856/0.9581 0.1517/0.9816 0.1930/0.9744 0.1318/0.9817 13.09
henrys law constant for gas in water 0.5454/0.3000 0.3017/0.7939 0.1773/0.9430 0.3234/0.6547 -82.37
Table 7: Results of thermal properties.
Property Unit Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
melting point K 19.9979/0.8199 14.1449/0.9053 24.7538/0.8050 11.0789/0.9171 21.68
boiling point K 19.8744/0.8891 6.2177/0.9906 12.3314/0.9777 4.3259/0.9933 30.43
vapor pressure mmHg 0.4837/0.5674 0.1596/0.9579 0.2000/0.9370 0.1365/0.9605 14.47
Table 8: Results of structural properties.
Property Unit Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
density of liquid g/ml 0.0331/0.9544 0.0160/0.9761 0.0145/0.9763 0.0120/0.9622 17.03
acentric factor omega 0.0702/0.7591 0.0233/0.9293 0.0225/0.9312 0.0169/0.9043 24.71
refractive index - 0.0277/0.4260 0.0057/0.9314 0.0079/0.9276 0.0046/0.9529 20.18
radius of gyration $10^{-10}$m 0.3573/0.6317 0.1230/0.9715 0.1204/0.9757 0.1085/0.9803 9.92
dipole moment D 0.4022/0.6638 0.2987/0.8115 0.3256/0.7585 0.3326/0.7610 -11.36
Table 9: Results of energetic properties.
Property Unit Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
entropy of formation J/mol/K 90.1219/0.8775 16.6006/0.9975 11.1386/0.9982 7.5413/0.9983 32.30
gibbs energy of formation J/mol 32.5168/0.9365 8.3799/0.9865 19.1059/0.9785 4.6748/0.9972 44.21
helmholtz energy of formation kJ/mol 30.2883/0.9286 7.9805/0.9829 24.8331/0.9656 5.4749/0.9746 31.40
internal energy of formation kJ/mol 34.2535/0.8869 10.3932/0.9795 36.3480/0.9393 5.7416/0.9950 44.76
enthalpy of vaporization kJ/mol 3.3764/0.7398 1.3875/0.9694 1.2207/0.9773 1.0592/0.9759 13.23
enthalpy of fusion J/mol/K 3.3815/0.8819 1.2265/0.9898 1.3044/0.9888 0.8394/0.9899 31.56
enthalpy of combustion kJ/mol 664.8247/0.8814 62.4789/0.9986 39.3110/0.9987 15.1036/0.9994 61.58
entropy of gas J/mol/K 38.1690/0.8946 9.9913/0.9900 10.9098/0.9896 7.2848/0.9898 27.09
enthalpy of formation kJ/mol 43.0063/0.8548 12.9888/0.9376 12.3369/0.9486 6.6573/0.9648 46.04
Table 10: Results of transport properties.
Property Unit Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
viscosity of liquid mp 0.8216/0.5870 0.4716/0.8328 0.9996/0.5924 0.3192/0.8905 32.32
thermal conductivity of gas W/m/K 0.0019/0.2592 0.0003/0.9584 0.0006/0.9293 0.0002/0.9572 20.00
thermal conductivity of liquid W/m/K 0.0107/0.0565 0.0028/0.7911 0.0036/0.8183 0.0027/0.7895 4.29
diffusion coefficient at infinite dilution in water cm2/s 0.0000/-1.2725 0.0000/0.9865 0.0000/0.9817 0.0000/0.9904 -
diffusion coefficient in air cm2/s 0.0066/0.6975 0.0014/0.9845 0.0011/0.9858 0.0009/0.9878 17.27
Table 11: Results of other properties.
Property Unit Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
surface tension dynes/cm 1.5226/0.8367 1.0414/0.8951 1.7084/0.8563 0.8205/0.9218 21.21
hydration free energy kJ/mol 0.8510/0.8748 0.4545/0.9646 0.4627/0.9623 0.4521/0.9630 0.52
Critical and Saturation Properties
Suiren-ConfAvg attains the lowest MAE across all five critical properties, with relative improvements ranging from 19.70% (critical compressibility) to 48.88% (critical volume) over the strongest baseline. Notably, while Uni-Mol v1 achieves marginally higher R² on critical temperature and density, Suiren-ConfAvg maintains competitive R² values (>0.97) while substantially reducing prediction errors, indicating superior calibration for extreme-value regression tasks.
Safety and Fluctuation Properties
For safety-related properties, Suiren-ConfAvg consistently outperforms baselines in both MAE and R², with the most pronounced gain observed for upper explosive limit (22.07% MAE reduction). In fluctuation properties, the method demonstrates exceptional capability in modeling solid-phase heat capacity (67.40% improvement).
Solution Properties
Suiren-ConfAvg achieves best-in-class performance on five of six solution properties. Improvements are particularly notable for solubility prediction in pure and saline water (15.07% and -10.07% MAE, respectively), which are critical for pharmaceutical and environmental applications. However, for Henry’s law constant of gases in water, the method underperforms Uni-Mol v2 by a substantial margin (−82.37%). We hypothesize this stems from the sparse and heterogeneous distribution of gas-phase solubility data, which may require specialized augmentation strategies.
Thermal and Structural Properties
Across thermal properties, Suiren-ConfAvg reduces MAE by 14.47%–30.43% while maintaining R² ≥ 0.947. For structural descriptors, the method excels in predicting liquid volume (41.83% improvement) and surface tension (21.21%), reflecting its capacity to encode intermolecular interaction patterns. The sole exception is dipole moment, where Uni-Mol v1 retains a edge (MAE: 0.2987 vs. 0.3326, with Suiren underperforming by 11.36%).
Energetic and Transport Properties
The most consistent gains are observed in energetic properties, where Suiren-ConfAvg achieves optimal MAE on all nine tasks, with improvements of 44.21% for Gibbs energy, 44.76% for internal energy, and 46.04% for enthalpy of formation. This suggests that the energy-related knowledge learned by Suiren-Base is transferred to Suiren-ConfAvg. For transport properties, substantial improvements are seen in viscosity (32.32%) and gas-phase thermal conductivity (20.00%), though liquid-phase thermal conductivity shows modest gain (4.29%), possibly due to stronger dependence on many-body hydrodynamic effects.
5.2.2 TDC ADMET group
We evaluated the performance of Suiren-ConfAvg on the TDC ADMET benchmarks. All experiments strictly adhered to the official evaluation protocols and metric settings provided by TDC to ensure a fair comparison. The results for regression tasks (MAE), classification tasks (AUROC and AUPRC) are presented in Tables 12, 13, and 14, respectively.
A critical distinction of Suiren-ConfAvg lies in its training protocol. Unlike several methods that may rely on extensive task-specific hyperparameter optimization, Suiren-ConfAvg was evaluated using a single, fixed training configuration across all datasets without complex hyperparameter search.
Table 12: Comparison of results for regression properties in TDC ADMET (MAE).
Property ChemProp DeepAutoQSAR DeepPurpose Uni-QSAR Suiren-ConfAvg
Caco2 0.3900 0.3060 0.3930 0.2730 0.3037
Lipophilicity 0.4370 0.4760 0.5740 0.4200 0.3865
AqSol 0.8200 0.7840 0.8270 0.6770 0.6794
PPBR 7.9930 8.0430 9.9940 7.5300 8.3963
LD50 0.5480 0.5900 0.6780 0.5530 0.5317
Table 13: Comparison of results for classification properties in TDC ADMET (AUROC).
Property ChemProp DeepAutoQSAR DeepPurpose Uni-QSAR Suiren-ConfAvg
HIA 0.9790 0.9820 0.9720 0.9920 0.9963
Pgp 0.9020 0.9170 0.9180 0.9340 0.9351
Bioavailability 0.6230 0.6820 0.6720 0.7320 0.7363
BBB 0.8820 0.8760 0.8890 0.9250 0.9236
CYP3A4 Substrate 0.6100 0.6420 0.6390 0.6450 0.6286
hERG 0.7500 0.8450 0.8410 0.8560 0.8816
Ames 0.8640 0.8640 0.8230 0.8760 0.8433
DILI 0.9180 0.9330 0.8750 0.9420 0.9378
Table 14: Comparison of results for classification properties in TDC ADMET (AUPRC).
Property ChemProp DeepAutoQSAR DeepPurpose Uni-QSAR Suiren-ConfAvg
CYP2C9 Inhibition 0.7700 0.7920 0.7420 0.8010 0.7845
CYP2D6 Inhibition 0.6640 0.7020 0.6160 0.7430 0.6725
CYP3A4 Inhibition 0.8700 0.8830 0.8290 0.8880 0.8740
CYP2C9 Substrate 0.3910 0.3950 0.3800 0.4540 0.5177
CYP2D6 Substrate 0.6880 0.7030 0.6770 0.7210 0.7289
Despite this constraint, the model achieved SOTA results in 8 out of 18 total metrics and ranked second in an additional 4 metrics. In cases where Suiren-ConfAvg did not secure the first place (e.g., AqSol, BBB, CYP3A4 Substrate), the performance gaps were negligible. This suggests that the performance sacrifice, if any, is minimal compared to the gains in reproducibility and ease of deployment. The ability to deliver highly competitive, often leading, performance across regression and classification tasks without fine-tuning highlights the strong generalization capability and robustness of the Suiren-ConfAvg architecture. These results validate that Suiren-ConfAvg offers an efficient and reliable solution for ADMET prediction, balancing high predictive accuracy with practical implementation simplicity.
6 Conclusion
In this work, we propose the Suiren-1.0 family, which comprises three models: Suiren-Base, Suiren-Dimer, and Suiren-ConfAvg. Suiren-Base and Suiren-Dimer are two 3D conformational models, whose performance is ensured through large-scale pre-training. Suiren-ConfAvg is obtained by distilling the 3D representations from Suiren-Base into the 2D representation space via our proposed CCD method. We have validated the strong performance of Suiren-1.0 on various molecular tasks through extensive experiments. The models and benchmarks developed in this work have also been open-sourced. We hope this work can support research on molecular foundation models.
Suiren-1.0 also has some limitations and directions for future work: (1) Due to computational constraints, we are unable to further scale up the model size; (2) In the MoE framework of Suiren-Base, we adopted a dense expert strategy—in the future, as the number of experts increases, we can use Top-K to improve inference speed; (3) For some specific downstream tasks, the potential of Suiren models can be further explored through hyperparameter search.
Appendix
Appendix A Contributions and Acknowledgments
Research & Engineering & Data Computing
Junyi An
Xinyu Lu (Intern)
Yun-Fei Shi
Li-Cheng Xu
Nannan Zhang (Intern)
Chao Qu
Fenglei Cao
Yuan Qi
We would also like to acknowledge the SAIS platform and all members of Golab, who contributed to the development of the Suiren-1.0 model in critical areas such as business and evaluation operations.
Appendix B The explanation of evaluation metrics in Figure 1 and Figure 2
In Figure 1, we use the inverse MAE to quickly demonstrate the performance of Suiren across various properties. Specifically, the MAE values were normalized and mapped to a standardized scoring scale ranging from 60 to 100. Since MAE represents prediction error (where lower values indicate better performance), an inverse Min-Max normalization strategy was employed.
For a specific property $p$, let $E_{m,p}$ denote the MAE of model $m$. The normalized score $S_{m,p}$ for model $m$ on property $p$ is calculated as follows:
$S_{m,p}=60+40\times\frac{\max(E_{p})-E_{m,p}}{\max(E_{p})-\min(E_{p})+\varepsilon}$ (14)
where:
  • $\max(E_{p})$ and $\min(E_{p})$ represent the maximum and minimum MAE values observed among all evaluated models for property $p$, respectively.
  • The constant $60$ establishes the baseline score for the worst-performing model (i.e., when $E_{m,p}=\max(E_{p})$).
  • The scaling factor $40$ expands the score up to a maximum of $100$ for the best-performing model (i.e., when $E_{m,p}=\min(E_{p})$).
  • $\varepsilon$ is an infinitesimally small constant (e.g., $10^{-10}$) added to the denominator to prevent potential division-by-zero errors in cases where all models yield identical MAE values.
This transformation ensures that properties with drastically different magnitude scales are projected onto a uniform visual space, allowing for a fair, area-based geometric comparison in the resulting radar charts.
In Figure 2, we use scaled MAE. For each property, we divide the results of the four models by the maximum MAE value among the four models. This way, all results are unified to the range between 0 and 1. This approach alleviates the loss of visual information caused by large numerical differences between different properties.
Appendix C Evaluation of MoleHB Size-Stratified split
To further explore the generalization capability of foundation models under distribution shift, we systematically evaluated MoleBERT, Uni-Mol v1, Uni-Mol v2, and Suiren-ConfAvg on the Size-Stratified split subset of MoleHB, a setting that rigorously tests extrapolation to unseen molecular sizes. The comprehensive results across eight property categories are summarized in Tables 1522.
Overall performance trends
As anticipated, Size-Stratified split induces a non-trivial distribution shift, leading to performance degradation across all methods relative to random-split evaluations. Nevertheless, Suiren-ConfAvg demonstrates superior robustness: it achieves the lowest mean absolute error (MAE) on properties across all categories, with relative improvements ranging from 4.68% (critical pressure) to 91.20% (Helmholtz energy of formation) over the strongest baseline. Notably these confidence-weighted ensemble approaches achieve substantial improvements particularly on energetic and critical properties, underscoring the effectiveness of confidence-weighted ensemble strategies in mitigating size-induced generalization gaps.
Table 15: Results of critical and saturation properties: model performance (MAE/R2). Best MAE and best R2 per property are boldfaced.
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Critical Compressibility 0.03447 0.03632 0.03527 0.00907 73.69
Critical Density 0.00671 0.00720 0.00734 0.00888 -23.33
Critical Pressure 1.73699 3.27665 2.63610 1.65567 4.68
Critical Temperature 77.96931 63.34677 87.88906 19.59443 69.07
Critical Volume 351.31897 290.12731 321.89835 55.80474 80.77
Table 16: Results of safety properties.
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Flash Point 32.80089 20.89532 36.13503 14.83841 28.99
Lower Explosive Limit 0.10239 0.05315 0.10274 0.03367 36.65
Upper Explosive Limit 0.74490 0.59120 0.62222 0.34410 41.80
Table 17: Results of fluctuation properties.
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Heat Capacity Of Gas 144.17387 88.39811 146.68384 24.45830 72.33
Heat Capacity Of Liquid 60.96739 26.91383 40.76889 15.74570 41.50
Heat Capacity Of Solid 61.55533 43.78053 71.91361 4.97646 88.63
Coefficient Of Thermal Expansion Of Liquid 0.00010 0.00005 0.00009 0.00004 20.00
Table 18: Results of solution properties.
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Octanol Water Partition Coefficient 0.68611 0.29471 0.44365 0.19774 32.90
Solubility In Water 0.32445 0.08128 0.28590 0.07426 8.64
Solubility In Water Containing Salt 0.25057 0.09037 0.17057 0.06786 24.91
Solubility Parameter 1.36914 1.53168 1.41960 0.75884 44.58
Table 19: Results of thermal properties.
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Boiling Point 66.21946 53.40313 70.61806 21.18527 60.33
Melting Point 39.98979 42.88753 51.46680 40.24486 -0.00
Vapor Pressure 1.21508 0.89771 1.08963 0.47305 47.30
Table 20: Results of structural properties.
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Density Of Liquid 0.01929 0.01221 0.01210 0.00851 29.67
Acentric Factor 0.09174 0.07457 0.08191 0.06685 10.35
Radius Of Gyration 1.15782 0.53795 0.73237 0.22784 57.65
Refractive Index 0.01290 0.00859 0.00974 0.00673 21.65
Dipole Moment 0.64795 0.49490 0.53720 0.57567 -7.16
Table 21: Results of energetic properties.
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Gibbs Energy Of Formation 99.86093 52.53548 74.17537 6.85050 86.96
Helmholtz Energy Of Formation 109.97257 73.68592 86.25709 6.48666 91.20
Internal Energy Of Formation 149.43255 90.63668 135.46958 11.44695 87.37
Enthalpy Of Combustion 4411.49423 3173.83835 3606.45797 338.80971 89.32
Enthalpy Of Formation 186.20484 123.64866 135.32938 11.15240 90.98
Enthalpy Of Fusion 13.58065 14.34938 15.46400 2.60221 80.84
Enthalpy Of Vaporization 9.32412 3.97332 6.68632 2.41336 39.26
Entropy Of Formation 543.36403 606.29039 688.57106 176.35892 67.54
Entropy Of Gas 258.35003 247.96703 218.72920 18.00675 91.77
Table 22: Results of transport properties.
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Thermal Conductivity Of Gas 0.00227 0.00139 0.00148 0.00072 48.20
Thermal Conductivity Of Liquid 0.00331 0.00276 0.00316 0.00203 26.45
Diffusion Coefficient At Infinite Dilution In Water 0.00000 0.00000 0.00000 0.00000 0.00
Diffusion Coefficient In Air 0.00487 0.00418 0.00363 0.00322 11.29
Viscosity Of Liquid 1.71584 1.40034 1.83852 1.27413 9.01
Category-wise analysis
  • Energetic properties (Table 21): This category exhibits the most pronounced performance disparity. Baseline methods suffer severe degradation (e.g., MoleBERT’s MAE for enthalpy of combustion is 4411.49), whereas Suiren-ConfAvg maintains substantially lower errors (338.81), representing an 89.32% relative improvement. We hypothesize that the pre-training objective of Suiren-Base, which incorporates physics-informed energy constraints, enables more transferable representations for thermodynamic quantities that are sensitive to subtle electronic and conformational features.
  • Critical and saturation properties (Table 15): Suiren-ConfAvg dominates in critical temperature (19.59 vs. 63.35, best baseline) and critical volume (55.80 vs. 290.13), with relative improvements of 69.07% and 80.77%, respectively. However, for critical density, MoleBERT achieves marginally better results (0.00671 vs. 0.00888). For critical pressure, Suiren achieves 1.66 vs. MoleBERT’s 1.74, with a 4.68% improvement. This suggests that certain intensive properties with strong linear correlations to molecular size may be adequately captured by simpler architectures, whereas extensive or composite properties benefit from Suiren’s enhanced representation capacity.
  • Thermal and fluctuation properties (Tables 1917): The density of liquid shows moderate performance variation across splits (Suiren-ConfAvg: 0.00851), consistent with its dependence on atomic composition, with a 29.67% improvement over the best baseline. In contrast, heat capacities exhibit substantial size sensitivity, with Suiren-ConfAvg achieving 41.50%–88.63% improvements for liquid and solid heat capacities respectively, while showing 72.33% improvement for gas heat capacity.
  • Structural properties (Table 20): While Suiren-ConfAvg excels in radius of gyration (57.65% improvement), it is slightly outperformed by Uni-Mol v1 on refractive index (0.00859 vs. 0.00673) and dipole moment (0.49490 vs. 0.57567 for Suiren).
  • Safety, solution, and transport properties (Tables 1622): Suiren-ConfAvg consistently attains the best or near-best performance, with particularly notable gains in flash point (28.99% improvement) and vapor pressure (47.30% improvement).
Implications for molecular foundation models
The pronounced robustness of Suiren-ConfAvg on energetically complex and size-sensitive properties suggests that integrating physics-aware pre-training objectives with uncertainty-aware inference mechanisms can substantially improve out-of-distribution generalization.
Appendix D Ablation Studies of Suiren-ConfAvg
To systematically assess the contributions of Suiren-ConfAvg and DGNN to downstream task performance, we designed a comprehensive ablation study with four distinct model variants: (1) Baselines: the complete architecture with DGNN+Suiren-ConfAvg serving as our baseline; (2) Suiren-ConfAvg only ( ): the DGNN component removed, with Suiren-ConfAvg acting as the sole backbone—prediction heads are attached to frozen Suiren-ConfAvg embeddings, allowing only the heads to be trained; (3) Suiren-ConfAvg only ( ): built upon variant (2) but with Suiren-ConfAvg weights unfrozen and co-optimized end-to-end; (4) w/o. Suiren-ConfAvg: Suiren-ConfAvg pre-trained weights entirely discarded, replaced by a deep task-specific GNN trained from scratch. To ensure the generalizability of our experimental findings, we selected properties spanning diverse domains to highlight the representational differences across distinct tasks.
Table 23: Results of ablation studies on MoleHB random split.
Property Baselines Suiren-ConfAvg Suiren-ConfAvg w/o.
only ( ) only ( ) Suiren-ConfAvg
Critical Temperature 6.6725 8.6190 7.3257 7.9764
Flash Point 8.2072 9.5567 8.8750 9.2798
Melting Point 11.0789 12.7397 12.4325 12.7382
Surface Tension 0.8205 0.9020 0.8568 0.9123
From the results presented in Table 23, we draw the following conclusions: (1) Necessity of task-specific adaptation: The removal of the task-specific GNN from the DGNN architecture precipitates significant performance degradation, with this effect being particularly pronounced when Suiren-ConfAvg weights remain frozen. This finding underscores that effective transfer of Suiren-ConfAvg representations to downstream tasks necessitates sufficient learnable capacity and modular adaptation, while simultaneously corroborating the existence of substantial distributional divergence across macroscopic property prediction tasks. (2) Efficacy of pre-trained representations: The exclusion of Suiren-ConfAvg pre-trained weights uniformly impairs performance across all evaluated tasks, yielding inferior results compared to the Suiren-ConfAvg-with-prediction-head configuration. This establishes the beneficial role of Suiren-ConfAvg representations in enhancing downstream predictive capabilities.
Interestingly, the task-specific GNN trained from scratch achieves competitive performance relative to prior molecular foundation models. This phenomenon may be attributed to two factors: (1) the inherent limitations of conventional molecular representations in capturing the broad spectrum of macroscopic physicochemical properties; (2) architectural innovations in our 2D GNN design, which integrates fully-connected GAT layers with adjacency-matrix-based GAT layers to enhance expressive capacity. These ablation studies provide compelling empirical evidence for the indispensability of synergistically combining Suiren-ConfAvg pre-trained representations with the adaptive capacity of the DGNN architecture.